# Generative Models From and For Sampling-Based MPC: A Bootstrapped Approach For Adaptive Contact-Rich Manipulation

### v1

@@ I Introduction
Reactive contact-rich (loco-)manipulation in high-dimensional state and action spaces poses significant challenges for real-time control. Sampling-based Model Predictive Control, or sampling-based predictive control (SPC), offers a principled framework to address these challenges by solving trajectory optimization problems online with a model in the loop, enabling adaptive behavior and constraint satisfaction [1, 2, 3, 4]. However, the computational cost of forward simulation, combined with the challenge of effectively exploring the search space in high-dimensional, contact-rich environments, limits the applicability of real-time optimization to more complex behaviors and higher-frequency control.
To address this, a promising line of work seeks to amortize the computational burden of online optimization by shifting it to an offline phase [5, 6, 7]. The key idea is to collect data, either from expert demonstrations or from model-based planning, and use it to train a generative model that captures the distribution of useful actions or control sequences. At test time, this model can then be used to guide or warmstart the sampling distribution in MPC, drastically improving efficiency and solution quality by focusing sampling on high-likelihood, constraint-satisfying regions of the action space.
Recent advances in generative modeling, particularly diffusion and flow-matching models, have shown strong performance in learning expressive policies for dexterous manipulation tasks [8, 9, 10]. Similarly, offline model-based reinforcement learning methods [11, 12] leverage precomputed data to enable fast runtime control via policy networks trained to approximate optimal solutions. Yet, these methods are often limited by the scope of their training data and struggle to generalize to out-of-distribution (OOD) states or tasks. In response to these limitations, several recent works demonstrate that bootstrapping online planners with offline-trained generative models leads to faster convergence, better exploration, and more robust performance in complex environments [7, 13, 14, 15]. In this paper, we adopt this amortized optimization perspective, focusing on how offline data collection and generative modeling can be used to accelerate and guide online sampling-based MPC in contact-rich, high-dimensional settings while maintaining the flexibility and adaptability of online optimization.
@@
We propose a generative predictive control (GPC) framework that bootstraps SPC with conditional flow-matching models trained on SPC control sequences collected in simulation. To the best of our knowledge, we are the first to show that meaningful proposal distributions can be learned directly from noisy SPC data without requiring expert refinement or numerical solvers. We are also the first to demonstrate this approach on real hardware in a contact-rich loco-manipulation task, and show that it improves sample efficiency and generalizes robustly to task variations in both simulation and real world.
Figure 1: Generative predictive control (GPC) framework for bootstrapping sampling-based Model Predictive Control (SPC). We collect open-loop control sequences from an SPC algorithm in simulation and use them to train a generative proposal distribution. At test time, this model guides and amortizes online MPC, enabling non-myopic, constraint-satisfying behavior with improved sample efficiency and robustness in contact-rich, high-dimensional settings.
@@ II Related Work
SPC has been widely adopted for its robustness to nonlinearity, discontinuities, and nonconvexities, particularly in contact-rich robotic tasks [16, 17, 2, 1, 3, 18, 19]. These methods typically optimize a trajectory distribution by iteratively sampling candidate controls and selecting elites based on forward-simulated costs. However, their performance is often limited by the computational cost of forward dynamics, especially when simulating contact interactions or systems with many degrees of freedom (DOF). While prior work has sought to speed up forward simulation, e.g., via quasi-static approximations [20] or learned dynamics models [21], these approaches often trade off fidelity or depend on highly accurate model learning. In contrast, we do not aim to replace the dynamics model but rather to amortize the trajectory optimization itself. We do this by learning generative models over control distributions derived from open-loop control sequences that either led to task success or incurred low cost in offline sampling-based MPC, enabling faster and more informed sampling at test time.
@@
Recent work leverages offline data to reduce the computational burden of of control algorithms at runtime. Learning from expert demonstrations or planner rollouts, via behavior cloning [8, 9, 10] or model-based RL [11, 12], has become a common approach. When sourced from high-quality planners, such as sampling-based MPC, this data enables training generative models that can guide or initialize online control. Several works have explored learning from planner-generated trajectories to approximate optimal or near-optimal solutions and accelerate planning [22, 23, 6, 24]. This idea underpins Approximate MPC (AMPC), where learned models bootstrap or replace expensive solvers. Early efforts used regression or Gaussian Mixture Models to represent multi-modal action distributions [25, 26], while recent work adopts diffusion and flow-matching models for their expressiveness [27, 7, 28]. For instance, [29] use diffusion models to approximate near-globally optimal MPC solutions from locally optimal trajectories generated by a numerical solver [30]. At runtime, a sample-score-rank strategy is applied to select the best candidate sampled from the learned distribution. While conceptually similar to our approach, their results are limited to simple environments due to the slow inference speed of the diffusion model. Our work is most closely related to Kurtz et al. [14], who also boostrap SPC using generative models trained on SPC control sequences. Their method uses online data, performing iterative training that alternates between data collection and model updates, using controls sampled from each partially trained model. While this strategy aims to improve the quality of the training data, we find empirically that it can, in fact, degrade performance by collapsing the multi-modal structure of the control sequence distribution, crucial for effective sampling. In contrast, our method trains a generative model over offline data (noisy SPC rollouts) and achieves strong performance, showing that bootstrapped SPC does not require iterative refinement. To our knowledge, [14] and our work are the first to learn from SPC data, rather than from trajectories generated by gradient-based solvers.
@@
Another line of work aims to improve SPC by learning structured priors over control sequences in latent action spaces using generative models [31, 32]. These methods typically rely on expert demonstrations, which lack exploratory diversity, and require complex bi-level training to learn both latent spaces and their distributions. In contrast, we focus on directly leveraging the diverse data that can be produced by sampling-based MPC in an offline data collection setting. Freed from real-time constraints, we can expand the search space during planning, yielding richer control sequences that support efficient sampling through simpler generative models at test time.
@@ III-A Problem Formulation
In this paper we consider optimal control problems in continuous action spaces. Given an initial state $\bm{x}_{0}=\bm{x}_{init}$, the objective is to determine a sequence of open-loop control actions $U_{\tau}=[\bm{u}_{\tau},\bm{u}_{\tau+1},\dots,\bm{u}_{\tau+T}]$ that minimizes a given cost function over a finite time horizon $T$:
$\min_{\bm{u}_{0},\bm{u}_{1},\ldots,\bm{u}_{T}}\quad$ $\phi(\bm{x}_{T+1})+\sum_{\tau=0}^{T}\ell(\bm{x}_{\tau},\bm{u}_{\tau})$ (1a)
@@
where $\bm{x}_{\tau}\in\mathbb{R}^{n}$ and $\bm{u}_{\tau}\in\mathbb{R}^{m}$ are the state vector and the control input at time step $\tau$, respectively. The functions $\ell:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ and $\phi:\mathbb{R}^{n}\to\mathbb{R}$ represent the stage and terminal cost, respectively. We do not assume to know the closed-form expression for the system dynamics $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$, but instead assume access to a simulator (e.g. MuJoCo [33]) or a learned model. For a more compact notation, we define a cost function $J:\mathbb{R}^{m\times T}\times\mathbb{R}^{n}\to\mathbb{R}$ that encapsulates both, costs and system dynamics, allowing us to write the problem as
@@ III-B Sampling-based MPC (SPC)
Contact-rich robot control tasks pose significant challenges due to the non-convexity of cost functions and the nonlinear, often discontinuous nature of system dynamics. Sampling-based MPC addresses these issues by optimizing over a parameterized distribution $\pi_{\phi}(U)$ rather than directly computing the optimal control sequence.
We consider a generic SPC procedure in which, at each control step $\tau$, the controller samples $N$ control sequences $\{U^{(i)}\}_{i=1}^{N}$ from the current distribution $\pi_{\phi}$, simulates their outcomes from the current state estimate $\hat{\bm{x}}_{\tau}$, and evaluates them using the cost function $J(U^{(i)};\hat{\bm{x}}_{\tau})$. Based on these evaluations, the distribution parameters $\phi$ are updated according to the specific SPC algorithm used. The executed control $\bm{u}_{\tau}$ is typically the first element of $\bar{U}_{\tau}$, or alternatively, it can be derived via spline interpolation across the optimized sequence. While other works explore non-Gaussian proposal distributions [31], we focus on diagonal Gaussian distributions of the form $\pi_{\phi}(U)=\mathcal{N}(\bar{U},\Sigma)$, as used in the Cross-Entropy Method (CEM) [34] and other SPC algorithms [2, 16]. Here, $\phi=(\bar{U},\Sigma)$, and $\Sigma=\mathrm{diag}(\bm{s})$, with $\bm{s}$ denoting a vector of variances across dimensions.
@@ III-C Generative Modeling: Flow-Matching
While the above focuses on shaping a sampling distribution for SPC, generative modeling focuses on a different problem: produce a sample $x$ from a target distribution $p(x)$, which is typically unknown in closed form, but can be approximated by a dataset of samples $\mathcal{D}$. Among recent approaches to generative modeling, two closely related approaches have gained significant traction due to their ability to capture complex, multi-modal distributions: flow matching [35] and diffusion models [36]. The underlying concept is to learn a distribution over trajectories or transformations that maps a simple prior distribution to complex target data. In this work, we focus on flow-matching models and their conditional variant [37], as they offer superior inference speed compared to diffusion models. Flow matching aims to learn a time-dependent vector field $v_{\theta}(\mathbf{x},t)$ that transports samples from an easy-to-sample prior distribution $p_{0}(\mathbf{x})$ (e.g., a standard Gaussian) to a target data distribution $p_{1}(\mathbf{x})$. The flow network $v_{\theta}$ is trained on
$\min_{\theta}\;\mathbb{E}_{\begin{subarray}{c}\mathbf{x}_{0}\sim p_{0},\\ \mathbf{x}_{1}\sim p_{1},\\ t\sim\mathcal{U}(0,1)\end{subarray}}\left[\left\|v_{\theta}\left(t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0},\,t\right)-(\mathbf{x}_{1}-\mathbf{x}_{0})\right\|^{2}\right]$ (1)
This loss encourages the vector field to push samples along straight-line paths from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$. At inference time, new samples are generated by drawing $\mathbf{x}_{0}\sim p_{0}(\mathbf{x})$ and solving the ODE $\dot{\mathbf{x}}=v_{\theta}(\mathbf{x},t)$ from $t=0$ to $t=1$, typically approximated with an Euler scheme $\mathbf{x}_{t+\delta t}=\mathbf{x}_{t}+v_{\theta}(\mathbf{x}_{t},t)\delta t$.
@@ IV Bootstrapping Sampling-based MPC with Generative Flow-Matching Models
In this section, we introduce our approach for bootstrapping SPC with generative models trained on open-loop control sequences collected from SPC itself. The core idea is to learn a generative model that approximates the distribution of successful control sequences conditioned on the current state and history. At test time, this model serves as a proposal distribution to guide and warm-start the sampling process in online SPC, improving sample efficiency and robustness.
@@ IV-A Data collection
The offline data collection phase is central to our approach, as it should provide control sequences, that are likely to lead to task success, conditioned on the current state and history of states. Since we aim to bootstrap SPC at test time, we generate this dataset directly from an SPC algorithm, specifically CEM as it is widely used and easy to tune for different tasks, without the runtime constraints of online control. This allows for longer horizons and larger sample sizes to collect high-quality, non-myopic control sequences.
Given a task and associated cost function, we run the SPC algorithm across multiple episodes, each with random state initializations and capped at a maximum number of MPC iterations. During each episode, we record tuples of the form $(\bm{x}_{\tau},\bm{h}_{\tau},U^{*}_{\tau})$, where $\bm{x}_{\tau}$ is the current state, $\bm{h}_{\tau}$ encodes a fixed-window history of states, and $U^{*}_{\tau}$ denotes denotes the mean control sequence of the CEM distribution at time $\tau$. An experiment is considered successful if the task is completed within the allowed time steps. We define $\mathcal{I}_{\text{success}}$ as the index set of all successful experiment episodes, and construct our training dataset as
@@
To reduce the complexity of the generative model and improve runtime efficiency, as well as smoothness, each control sequence $U_{\tau}$ is represented using $K<H$ control points over a planning horizon of $H$ time steps, i.e. $U=[\bm{u}_{0},\bm{u}_{1},\dots,\bm{u}_{K}]$. These control points may correspond to spline knots or waypoints for linear interpolation. We also employ i) a progress-based heuristic to reset the variances during CEM in order to avoid early mode collapse, and ii) action-level annealing [18] that increases exploration, i.e., variances, for control points further into the horizon.
@@ IV-B Learning Control Sequence Proposal Distributions
Once we have collected a dataset of open-loop trajectories for a given task, we can train a flow-matching generative model to learn a time-varying state-conditional vector field $v_{\theta}(U,\bm{x}_{\tau},\bm{h}_{\tau},t)$ that pushes samples from $U_{t}\sim\mathcal{N}(0,I)$ at flow step $t=0$ to the target distribution $U_{t=1}\sim p_{\theta}(U\mid\bm{x}_{\tau},\bm{h}_{\tau})$ at $t=1$, i.e. the distribution of control sequences that are likely to lead to successful task completion given the current state $\bm{x}_{\tau}$ and state history $\bm{h}_{\tau}$. For simplicity, we summarize samples from the generative model as samples from the distribution $p_{\theta}(U\mid\bm{x}_{\tau},\bm{h}_{\tau})$. We refer to our method as generative predictive control (GPC), which leverages a learned distribution over control sequences conditioned on task context. This distribution can be used in two distinct ways: i) as a sampling distribution in a random shooting approach, where we sample a batch of control sequences from $p_{\theta}$ and select the best candidate based on the rollouts and their respective costs, or ii) in conjunction with an SPC algorithm, where samples from $p_{\theta}(U\mid\bm{x}_{\tau})$ are used to update the sampling distribution of the SPC algorithm, e.g., $\pi_{\phi}(U)$ in the CEM. In the following, we call the first approach GPC-Shoot and the second approach GPC-CEM.

### v2

@@ I Introduction
Reactive contact-rich (loco-)manipulation in high-dimensional state and action spaces poses significant challenges for real-time control. Sampling-based Model Predictive Control (SPC) offers a principled framework to address these challenges by solving trajectory optimization problems online with a model in the loop, enabling adaptive behavior and constraint satisfaction [24, 15, 19, 2]. However, the computational cost of forward simulation, combined with the challenge of effectively exploring the search space in high-dimensional, contact-rich environments, limits the applicability of real-time optimization to more complex behaviors and higher-frequency control.
A promising line of work seeks to amortize the computational burden of online optimization by shifting it to an offline phase [17, 10, 41]. The key idea is to collect high quality data and train a generative model to capture a distribution of useful actions or control sequences. At test time, this model can be used to guide or warmstart the sampling distribution. The method attempts to drastically improve solution quality and efficiency by focusing sampling on high-likelihood, constraint-satisfying regions of the action space.
Recent advances in generative modeling, particularly diffusion and flow-matching models, have shown strong performance in learning expressive end-to-end policies for dexterous manipulation tasks [3, 6, 40]. Offline model-based reinforcement learning methods [11, 7] also show strong performance approximating optimal solutions by leveraging precomputed data to enable fast runtime control via policy networks. However, these methods are often limited by the scope of their immense training data and struggle to generalize to out-of-distribution (OOD) states or tasks. In response to these limitations, several recent works demonstrate that bootstrapping online planners with offline-trained generative models leads to faster convergence, better exploration, and more robust performance in complex environments [41, 30, 22, 37]. In this paper, we focus on how offline data collection and generative modeling can both accelerate and guide online sampling-based MPC in contact-rich, high-dimensional settings while maintaining the flexibility and adaptability of online optimization.
@@
We propose a generative predictive control (GPC) framework that bootstraps SPC with conditional flow-matching models trained on SPC control sequences collected in simulation. To the best of our knowledge, we are the first to show that meaningful proposal distributions can be learned directly from noisy SPC data without requiring expert refinement or numerical solvers. We are also the first to demonstrate that this approach improves sample efficiency and generalizes robustly to task variations in both simulation andreal hardware in a contact-rich loco-manipulation task.
Figure 1: Generative predictive control (GPC) framework for bootstrapping sampling-based MPC. We collect open-loop control sequences from an SPC algorithm in simulation and use them to train a generative proposal distribution. At test time, this model guides and amortizes online MPC, enabling non-myopic, constraint-satisfying behavior with improved sample efficiency and robustness in contact-rich, high-dimensional settings.
@@ II Related Work
SPC has been widely adopted for its robustness to nonconvex and discontinuous problems, particularly in contact-rich robotic tasks [38, 21, 15, 24, 19, 39, 27]. These methods typically optimize a trajectory distribution by iteratively sampling candidate controls and selecting actions based on forward-simulated costs. Their performance is often limited by the computational cost of forward dynamics, especially when simulating contact interactions or systems with many degrees of freedom (DOF). While prior work has sought to speed up forward simulation, e.g., via quasi-static approximations [28] or learned dynamics models [18], these approaches often trade off fidelity or depend on highly accurate model learning. In contrast, we do not aim to replace the dynamics model but rather to amortize the trajectory optimization itself. We do this by learning generative models over control distributions derived from open-loop control sequences that either led to task success or incurred low cost in offline sampling-based MPC, enabling faster and more informed sampling at test time.
@@
Recent work leverages offline data to reduce the computational burden of control algorithms at runtime. Common approaches include behavior cloning on expert demonstrations and planner rollouts [3, 6, 40] or model-based RL [11, 7]. When sourced from high-quality planners (such as sampling-based MPC), this data enables training generative models that can guide or initialize online control. Several works have explored learning from planner-generated trajectories to approximate optimal solutions and accelerate planning [8, 5, 10, 35]. This idea underpins Approximate MPC (AMPC), where learned models bootstrap or replace expensive solvers [4]. A representative example [16] uses diffusion models to approximate near-globally optimal MPC solutions from locally optimal trajectories generated by a numerical solver [36]. Yet, the learned models are not used to guide sampling but rather to replace the solver entirely. Our work is most closely related to [22], which also bootstraps SPC using generative models trained on SPC control sequences. Their method alternates online data collection with model updates, but we find this iterative refinement can collapse the multi-modal control distribution important for effective sampling. In contrast, our method trains a generative model over offline data (noisy SPC rollouts) and achieves strong performance, showing that bootstrapped SPC does not require iterative refinement. To the best of our knowledge, both [22] and our approach are the first to learn from SPC data rather than trajectories from gradient-based solvers.
@@
Another line of work aims to improve SPC by learning structured priors over control sequences in latent action spaces using generative models [32, 29]. These methods typically rely on expert demonstrations, which lack exploratory diversity and require complex bi-level training to learn both the latent spaces and their distributions. In contrast, we focus on directly leveraging the diverse data produced by sampling-based MPC during offline data collection. Freed from real-time constraints, we can instead expand the search space during planning to yield richer control sequences that support efficient sampling through simpler generative models.
Infinite-Horizon Value Approximation and MPC
A complementary line of research seeks to reduce the myopia of finite-horizon MPC by learning infinite-horizon value functions and integrating them into the control loop [20, 14, 26, 12]. These methods approximate an infinite-horizon value signal over states, which is then used as a terminal cost or shaping function for MPC, thereby injecting long-horizon foresight into an otherwise short-horizon optimizer. These approaches share the same motivation as ours, i.e., mitigating the short-horizon bias of online optimization, but are orthogonal in how they inject long-term structure. Even with an accurate estimate of the infinite-horizon value function, MPC still requires an effective mechanism for searching the control space. In contrast, our method focuses directly on improving this search by learning generative models over successful control sequences, thereby guiding the sampling distribution used by SPC.
@@ III-A Problem Formulation
In this paper we consider optimal control problems in continuous action spaces. Given an initial state $\bm{x}_{0}=\bm{x}_{init}$, the objective is to determine a sequence of open-loop control actions $U_{\tau}=[\bm{u}_{\tau},\bm{u}_{\tau+1},\dots,\bm{u}_{\tau+T}]$ that minimizes a given cost function $\ell(\bm{x},\bm{u})$ over a finite time horizon $T$:
$\min_{\bm{u}_{0},\bm{u}_{1},\ldots,\bm{u}_{T}}\quad$ $L_{f}(\bm{x}_{T+1})+\sum_{\tau=0}^{T}\ell(\bm{x}_{\tau},\bm{u}_{\tau})$ (1a)
@@
where $\bm{x}_{\tau}\in\mathbb{R}^{n}$ and $\bm{u}_{\tau}\in\mathbb{R}^{m}$ are the state vector and the control input at time step $\tau$, respectively. The functions $\ell:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ and $L_{f}:\mathbb{R}^{n}\to\mathbb{R}$ represent the stage and terminal cost, respectively. We assume access to a simulator (e.g. MuJoCo [33]) or a learned model to approximate the system dynamics $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$. For a more compact notation, we define a cost function $J:\mathbb{R}^{m\times T}\times\mathbb{R}^{n}\to\mathbb{R}$ that encapsulates both, costs and system dynamics, allowing us to write the problem as
@@ III-B Sampling-based MPC (SPC)
Contact-rich robot control tasks pose significant challenges due to non-convex cost functions and the nonlinear, often discontinuous nature of system dynamics. Sampling-based MPC addresses these issues by optimizing over a parameterized distribution $\pi_{\phi}(U)$ rather than directly computing the optimal control sequence. We consider a generic SPC procedure in which, at each control step $\tau$, the controller samples $N$ control sequences $\{U^{(i)}\}_{i=1}^{N}$ from the current distribution $\pi_{\phi}$, simulates their outcomes from the current state estimate $\hat{\bm{x}}_{\tau}$, and evaluates them using the cost function $J(U^{(i)};\hat{\bm{x}}_{\tau})$. Based on these evaluations, the distribution parameters $\phi$ are updated according to the chosen SPC algorithm. The executed control $\bm{u}_{\tau}$ is typically the first element of the sampled control sequence ${U}_{\tau}$ or derived via spline interpolation across the optimized sequence. We focus on diagonal Gaussian distributions of the form $\pi_{\phi}(U)=\mathcal{N}(\bar{U},\Sigma)$, as used in the Cross-Entropy Method (CEM) [31] and other SPC algorithms [15, 38]. Here, $\phi=(\bar{U},\Sigma)$, and $\Sigma=\mathrm{diag}(\bm{s})$, with $\bm{s}$ denoting a vector of variances.
@@ III-C Generative Modeling: Flow-Matching
While the above focuses on shaping a sampling distribution for SPC, generative modeling focuses on a different problem: produce a sample $x$ from a target distribution $p(x)$, which is typically unknown in closed form, but can be approximated by a dataset of samples $\mathcal{D}$. Among recent approaches to generative modeling, two closely related approaches have gained significant traction due to their ability to capture complex, multi-modal distributions: flow matching [25] and diffusion models [13]. The underlying concept is to learn a distribution over trajectories or transformations that maps a simple prior distribution to complex target data. In this work, we focus on flow-matching models and their conditional variant [34], as they offer superior inference speed compared to diffusion models. Flow matching aims to learn a time-dependent vector field $v_{\theta}(\mathbf{x},t)$ that transports samples from an easy-to-sample prior distribution $p_{0}(\mathbf{x})$ (e.g., a standard Gaussian) to a target data distribution $p_{1}(\mathbf{x})$.
@@ IV Bootstrapping Sampling-based MPC with Generative Flow-Matching Models
In this section, we introduce our approach to bootstrapping SPC with generative models trained on open-loop control sequences collected from SPC itself. The core idea is to learn a generative model that approximates the distribution of successful control sequences conditioned on the current state and history. At test time, this model serves as a proposal distribution to guide and warm-start the sampling process in online SPC, improving sample efficiency and robustness.
@@ IV-A Data collection
The offline data collection phase is central to our approach. It should provide control sequences (conditioned on the current state and history of states) likely to lead to task success. Since we aim to bootstrap SPC at test time, we generate this dataset directly from an SPC algorithm — specifically CEM, as it is widely used and easy to tune for different tasks without the runtime constraints of online control. This allows for longer horizons and larger sample sizes to collect high-quality, non-myopic control sequences.
Given a task and associated cost function, we run CEM across multiple episodes, each with random state initializations and capped at a maximum number of MPC iterations. During each episode, we record $(\bm{x}_{\tau},\bm{h}_{\tau},U^{*}_{\tau})$, where $\bm{x}_{\tau}$ is the current state, $\bm{h}_{\tau}$ encodes a fixed-window history of states, and $U^{*}_{\tau}$ denotes the mean control sequence of the CEM distribution at time $\tau$. An experiment is considered successful if the task is completed within the allowed time steps. We define $\mathcal{I}_{\text{success}}$ as the index set of all successful experiment episodes and construct our training dataset as
@@
To reduce the complexity of the generative model while improving runtime efficiency and smoothness, each control sequence $U_{\tau}$ is represented using $K<H$ spline interpolation points over a planning horizon of $H$ time steps, i.e. $U=[\bm{u}_{0},\bm{u}_{1},\dots,\bm{u}_{K}]$. We also employ i) a progress-based heuristic to reset the variances during CEM to avoid early mode collapse, and ii) action-level annealing [39] that increases exploration, i.e., variances, for control points further into the horizon.
@@ IV-B Learning Control Sequence Proposal Distributions
Once we have collected a task dataset of open-loop trajectories, we can train a flow-matching generative model to learn a time-varying state-conditional vector field $v_{\theta}(U,\bm{x}_{\tau},\bm{h}_{\tau},t)$ that pushes samples from the noise distribution $U_{t=0}\sim\mathcal{N}(0,I)$ to the target distribution $U_{t=1}\sim p_{\theta}(U\mid\bm{x}_{\tau},\bm{h}_{\tau})$, i.e. the distribution of control sequences that are likely to lead to successful task completion given the current state $\bm{x}_{\tau}$ and state history $\bm{h}_{\tau}$. For simplicity, we describe sampling from the generative model as sampling from the distribution $p_{\theta}(U\mid\bm{x}_{\tau},\bm{h}_{\tau})$. We refer to our method as generative predictive control (GPC), which leverages a learned distribution over control sequences conditioned on the task context. This distribution can be used in two distinct ways: i) to sample control sequences from using a random shooting approach and evaluate the best based on value functions, or ii) to update the sampling distribution of the SPC algorithm (e.g., $\pi_{\phi}(U)$ in CEM) with samples from $p_{\theta}(U\mid\bm{x}_{\tau},\bm{h}_{\tau})$. We call the first approach GPC-Shoot and the second approach GPC-CEM.